# Soft and Compliant Contact-Rich Hair Manipulation and Care

### v1

@@ Abstract
Hair care robots can help address labor shortages in elderly care while enabling those with limited mobility to maintain their hair-related identity. We present MOE-Hair, a soft robot system that performs three hair-care tasks: head patting, finger combing, and hair grasping. The system features a tendon-driven soft robot end-effector (MOE) with a wrist-mounted RGBD camera, leveraging both mechanical compliance for safety and visual force sensing through deformation. In testing with a force-sensorized mannequin head, MOE achieved comparable hair-grasping effectiveness while applying significantly less force than rigid grippers. Our novel force estimation method combines visual deformation data and tendon tensions from actuators to infer applied forces, reducing sensing errors by up to 60.1% and 20.3% compared to actuator current load-only and depth image-only baselines, respectively. A user study with 12 participants demonstrated statistically significant preferences for MOE-Hair over a baseline system in terms of comfort, effectiveness, and appropriate force application. These results demonstrate the unique advantages of soft robots in contact-rich hair-care tasks, while highlighting the importance of precise force control despite the inherent compliance of the system. Videos, data, and code are available at https://moehair.github.io.
Index Terms:
Assistive robotics, Soft robotics, Manipulation
@@ I Introduction
Hair plays an important role in people’s identities and self-esteem [1, 2]. Notably, the importance of hair to a person’s self-esteem tends to increase with age [3]. With aging and loss of independent mobility, hair care becomes an increasingly time-consuming and difficult daily task that is also crucial for personal hygiene [4]. Despite this, most elder care and hospice facilities heavily rely on volunteers for hair-care assistance  [5]. Toward addressing the gap in hair-care services, researchers proposed deploying robot assistance for combing [6, 7]. A notable challenge in previous works was that human subjects tend to perceive rigid robots as being “rough” [6]. Hair care and manipulation tasks additionally pose a perception challenge for robot systems because hair can often occlude the underlying scalp, making consistent application of force on the head difficult from just external vision. Previous approaches with mechanically rigid robot manipulators used high-cost force sensors to interact with the head safely [7].
@@
Soft robot manipulators have many advantages in addressing challenges in close contact with human users. Their compliance makes them safer in unstructured environments [8] and more robust in contact-rich manipulation tasks [9]. As such, soft robot manipulators are particularly useful in physical human-robot interaction (pHRI) tasks [10]. Human subjects tend to perceive soft robots as safer than their rigid counterparts [11]. Additionally, soft robot manipulators deform when they make contact, in contrast to rigid robots [12]. Observing these deformations offers a promising direction for using soft robots for interactive perception [13].
@@
Figure 1: Design of the proposed MOE end-effector. Left: exploded view and assembly of MOE. Right: fully assembled MOE.
@@
In this work, we propose a dexterous tendon-driven soft robot manipulator that we call Multi-fingered Omnidirectional End-effector (MOE) for hair-care applications (Fig. LABEL:fig:front). We first validate the physical safety of MOE using a force-sensorized mannequin head and then validate the efficacy of MOE for hair-care tasks in a user study. We demonstrate with a testbed that MOE’s compliant fingers make them appropriate for human contact. We show that MOE applies less force on the head compared to a rigid gripper, given the same hair depth, while being equally effective in grasping hair. Despite MOE’s compliance, the ability to reason about contact and applied forces is still important to maintaining contact with the user’s head and to apply consistent forces. To this end, we propose a method to infer MOE’s applied forces on the head by observing its deformation and the tendon tensions that are indirectly observed with the actuator current load. Finally, we present MOE-Hair, a system that incorporates the proposed MOE hand and the applied force estimation method to perform three hair care skills: head patting, finger combing and hair grasping. We find that users interacting with MOE rate it as effective, comfortable, and appropriately forceful across the three hair care skills we evaluated.
@@ II-A Hair Care Robots
Figure 2: MOE in varying poses. MOE deforms easily with contact with inherent passive compliance of the fingers’ material. As the fingers deform, they apply more tension on the tendons which our proposed method uses in conjunction with depth image to estimate applied forces. The tendons can be actuated in various ways to achieve compliant dexterity of the fingers.
@@ II-B Assistive Soft Robots
Researchers have studied soft robot manipulators’ advantages in various delicate manipulation tasks including food handling [8], crop handling [24], minimally invasive surgeries [25, 26, 27, 28] and wearable assistive robots [29]. In such applications, researchers noted that the inherent compliance of the soft robots enables reliable and safe performance [10].
@@
Figure 3: Setup for Training Data Collection. We use a controlled setup with a force sensor and allow MOE to randomly sample various contact conditions and capture training observations and forces in a self-labeling manner to scale up data collection.
@@ III-C Visual Perception Module
Assistive robot systems must be able to perceive and track the user’s pose to be effective and safe [16]. In this work, we are primarily concerned with tracking the user’s head pose. We deployed the Mediapipe perception model to extract the visual key points in the user’s face from the third-person RGBD camera’s RGB image and get the corresponding 3D points, which allows us to get head tracking results at 12.5 Hz. Fig. 6 shows the use of a visual marker on the tabletop to calibrate the coordinate frames among the third-person, wrist-mounted, and robot. The user’s 3D face key points are transformed into the robot frame for the subsequent hair manipulation tasks. For the finger combing task, we also used the Mediapipe model to extract key points on the participant’s hand to record the desired trajectory for MOE to track.
Figure 6: MOE-Hair System Setup. MOE is attached to a 6 degrees of freedom (DoF) robot arm. A third-person RGBD camera tracks key points of the head and the participant’s hand. A calibration marker is used to transform between camera and robot coordinate frames.
@@ III-D MOE-Hair System Integration
Figure 7: MOE-Hair System. We use a third-person RGBD camera to track the user’s visual face key points. The visual perception module maps the visual key points to the 3D point cloud and generates a task-specific trajectory. As MOE follows the trajectory, the system predicts the applied force on the head using the segmented depth images of the MOE fingers and actuator current load. The system adjusts the trajectories in real time to maintain comfort and contact based on the predicted contact forces.
@@ IV System Results
We present evaluation results for the components of the proposed MOE-Hair system. Specifically, we demonstrate that the proposed MOE soft robot end-effector can interact with a surface while applying lower contact forces compared to a rigid end-effector. Furthermore, we show that we can achieve improved applied force estimation using both observed depth and actuator current loads over using only depth or actuator load.
@@ IV-A MOE Interaction Force Evaluation
Figure 8: Visual comparison of the applied force estimation errors. We show a comparison between ground-truth forces experienced by the mannequin head with different wigs, the predicted forces from the baseline, and the proposed approaches.
@@ IV-B Applied Force Estimation Evaluation
For three wigs with different lengths and color, we have MOE follow a finger combing trajectory and record the depth image and actuator current load observation along with the synchronized ground-truth force readings from the sensor attached to the mannequin head. We visualize the root mean squared error (RMSE) over the trajectory in Fig. 8 and report them in Tab. II. In all three cases, the proposed MOE-Hair applied force estimation module outperforms both baselines, improving the actuator load-only and depth image-only baselines by up to 60.1 % and 20.3 % respectively.
@@ V User Study Evaluation
Figure 9: User study results evaluated with perceived task effectiveness, user comfort and appropriate use of force. Top row: results for the head patting task where MOE pats the user’s head from the top and the side. Middle row: results for the finger combing task where MOE follows the user’s demonstrated trajectory to comb the hair. Bottom row: results for the hair grasping task where MOE approaches the user’s head and grasps with its dexterous soft fingers.
@@ V-B Participants
We recruited 12 participants (ages 20–27, hair lengths 25–508 mm) without visual or mobility limitations as this is the first work with user studies on robot haircare. The participants represented diverse hair types based on an accepted classification [52]. The Institutional Review Board approved the protocol for the trials and the study.
@@ V-D Responses
Fig. 9 shows the participant responses for task effectiveness, user comfort, and the system’s appropriate use of force in each of the tasks for both force feedback conditions. Additionally, we report the participant responses for force feedback condition the participants preferred in Fig. 10, where higher Likert scale rating corresponds to a preference for MOE-Hair. We conducted non-parametric Wilcoxon signed-rank tests [54] for paired user responses based on the two systems performing the same tasks from Fig. 9 and also evaluated if the participants had a significant preference for a system from the neutral value of 4 with the responses in Fig. 10.
[…]
@@ V-E Qualitative Results
Effectiveness. All 12 of the participants responded positively to the system’s ability to perform the three skills outlined in this work. The participants indicated that MOE effectively performed the tasks and P6 noted, “It was working reliably." P12 noted, “I didn’t think it was going to work on my braids but it actually did well." Similarly, P2 remarked “I’ve been told my hair is difficult because it’s so thick" and commented that MOE-Hair performed effectively, noting “It did pretty well." The responses underscored the system’s surprising ability to perform well with varying hair characteristics and conditions.
Negative responses to the system’s effectiveness are related primarily to the participant’s uncertainty about the capabilities of the system. 3 of the participants responded that they wished the system could perform more tasks such as haircutting and styling, where P11 commented “it would’ve been nice if it could cut my hair," so that “I don’t have to talk to [human hairstylists]."
User tactile sensation. All 12 of the participants noted that they generally felt safe. For example, P7 noted that they “felt really comfortable." 5 participants commented that MOE’s contact sensation was pleasant. P1 liked the sensation when MOE made contact, remarking “It was like a head massage." P2 noted that MOE fingers felt “really similar to [human] fingers" and that “I kind of forgot it was a robot arm for a minute." P11 stated that MOE “made me relax."
Negative responses highlighted some participants’ hesitancy with an autonomous robot system touching their heads. P6 noted that MOE-Hair “felt like it’s kind of pushing my head." P12 noted that they were at first uncertain if the system was “safe or not" with the first task but stated that they became “more comfortable" with the latter tasks, suggesting the process of MOE-Hair system building trust with its users.
Appearance of MOE. 8 participants made generally positive comments about MOE’s appearance. As P4 noted, “I felt safe because it looked rubbery and soft." 3 participants noted that they liked that MOE did not look anthropomorphic and P5 stated MOE “just looks like it’s its own thing and it doesn’t give me an uncanny valley feeling."
2 participants noted that the wires of the actuators contributed to their unease with the system. For example, P1 noted MOE “seemed like a prototype and doesn’t seem like a product" because of the visible wires. P7 noted MOE “doesn’t look terrible" but that “the cables could be cleaned up." P12 made comments on MOE’s color and shape commenting “it doesn’t look friendly." Diverging comments on the appearance and visual behavior of the robot could indicate a need to incorporate design insights from literature [55] to be more likely to be perceived as being friendly by the users.
@@ VI Discussion and Limitations
In this work, we introduce a soft robot manipulator for hair manipulation and care tasks that we call MOE. The results suggest that in comparison to its mechanically rigid counterparts, MOE is safer in close contact with a head and that observing deformations of MOE can provide us with useful information about the applied forces on human users. The experimental results with the mannequin testbed showed that soft robots such as MOE could be effectively exploited in hair care tasks. Human user study highlighted that the proposed MOE-Hair system with the contact force estimation module was perceived favorably in its effectiveness for the tasks, comfort and ability to use appropriate forces. Qualitative analysis of the participants’ responses to open-ended questions revealed that some also perceived the tactile sensation of MOE to be pleasant and relaxing, which seems to offer promising extension in applying MOE to contact-rich pHRI tasks.
@@
Another limitation is in the system’s behavior generation. The focus of the MOE-Hair system development was on effective and safe contact using the compliance of the MOE soft fingers and force feedback with the proposed applied force estimation module. However, the characteristics of movement and personalization of robot trajectories could be important factors for users viewing the system favorably [6]. To this end, we can improve MOE-Hair system to incorporate strategies for planning trajectories with human preferences. Future work will also explore variable stiffness control with PCC to enhance force estimation accuracy and develop customized force thresholds and approach speeds for elderly and mobility-impaired users to maximize safety and comfort.

### v2

@@ Abstract
Hair care robots can help address labor shortages in elderly care while enabling those with limited mobility to maintain their hair-related identity. We present MOE-Hair, a soft robot system that performs three hair-care tasks: head patting, finger combing, and hair grasping. The system features a tendon-driven soft robot end-effector (MOE) with a wrist-mounted RGBD camera, leveraging both mechanical compliance for safety and visual force sensing through deformation. In testing with a force-sensorized mannequin head, MOE achieved comparable hair-grasping effectiveness while applying significantly less force than rigid grippers. Our novel force estimation method combines visual deformation data and tendon tensions from actuators to infer applied forces, reducing sensing errors by up to 60.1% and 20.3% compared to actuator current load-only and depth image-only baselines, respectively. A user study with 12 participants demonstrated statistically significant preferences for MOE-Hair over a baseline system in terms of comfort, effectiveness, and appropriate force application. These results demonstrate the unique advantages of soft robots in contact-rich hair-care tasks, while highlighting the importance of precise force control despite the inherent compliance of the system. Videos, data, and code are available at moehair.github.io.
@@ I Introduction
Hair plays an important role in people’s identities and self-esteem [1, 2, 3]. Notably, the importance of hair to a person’s self-esteem tends to increase with age [4]. With aging and loss of independent mobility, hair care becomes an increasingly time-consuming and difficult daily task that is also crucial for personal hygiene [5]. Despite this, most elder care and hospice facilities heavily rely on volunteers for hair-care assistance  [6]. Toward addressing the gap in hair-care services, researchers proposed deploying robot assistance for combing [7, 8]. A notable challenge in previous works was that human subjects tend to perceive rigid robots as being “rough” [7]. Hair care and manipulation tasks additionally pose a perception challenge for robot systems because hair can often occlude the underlying scalp, making consistent application of force on the head difficult from just external vision. Previous approaches with mechanically rigid robot manipulators used high-cost force sensors to interact with the head safely [8].
@@
Soft robot manipulators have many advantages in addressing challenges in close contact with human users. Their compliance makes them safer in unstructured environments [9] and more robust in contact-rich manipulation tasks [10]. As such, soft robot manipulators are particularly useful in physical human-robot interaction (pHRI) tasks [11, 12]. Human subjects tend to perceive soft robots as safer than their rigid counterparts [13]. Additionally, soft robot manipulators deform when they make contact, in contrast to rigid robots [14]. Observing these deformations offers a promising direction for using soft robots for interactive perception [15].
@@
Figure 1: Design of proposed MOE end-effector. Left: exploded view and assembly of MOE. Right: fully assembled MOE.
@@
In this work, we propose a dexterous tendon-driven soft robot manipulator that we call Multi-fingered Omnidirectional End-effector (MOE) for hair-care applications (Fig. LABEL:fig:front). We first validate the physical safety of MOE using a force-sensorized mannequin head and then validate the efficacy of MOE for hair-care tasks in a user study. We demonstrate with a testbed that MOE’s compliant fingers make them appropriate for human contact. We show that MOE applies less force on the head compared to a rigid gripper, given the same hair depth, while being equally effective in grasping hair. Despite MOE’s compliance, the ability to reason about contact and applied forces is still important to maintaining contact with the user’s head and to apply consistent forces. To this end, we propose a method to infer MOE’s applied forces on the head by observing its deformation and the tendon tensions that are indirectly observed with the actuator current load. Finally, we present MOE-Hair, a system that incorporates the proposed MOE hand and the applied force estimation method to perform three hair care skills: head patting, finger combing, and hair grasping. We find that users interacting with MOE rate it as effective, comfortable, and appropriately forceful across the three hair care skills we evaluated.
@@ II-A Hair Care Robots
Figure 2: MOE in varying poses. MOE deforms easily with contact by inherent passive compliance of the fingers’ material. As the fingers deform, they apply more tension on the tendons, which our proposed method uses in conjunction with depth image to estimate applied forces. The tendons can be actuated in various ways to achieve compliant dexterity of the fingers.
@@ II-B Assistive Soft Robots
Researchers have studied soft robot manipulators’ advantages in various delicate manipulation tasks including food handling [9], crop handling [26], minimally invasive surgeries [27, 28, 29, 30], and wearable assistive robots [31]. In such applications, researchers noted that the inherent compliance of the soft robots enables reliable and safe performance [11].
@@
Figure 3: Setup for training data collection. We use a controlled setup with a force sensor and allow MOE to randomly sample various contact conditions, capturing training observations and forces in a self-labeling manner to scale up data collection.
@@ III-C Visual Perception Module
Assistive robot systems must be able to perceive and track the user’s pose to be effective and safe [18]. In this work, we are primarily concerned with tracking the user’s head pose. We employ the MediaPipe perception model [52] to extract the visual key points in the user’s face from the third-person RGBD camera’s RGB image and get the corresponding 3D points, which allows us to get head tracking results at 12.5 Hz. Fig. 6 shows the use of a visual marker on the tabletop to calibrate the coordinate frames. The user’s 3D face key points are transformed into the robot frame for the subsequent hair manipulation tasks. For the finger combing task, we also use the MediaPipe model to extract key points on the participant’s hand to record the desired trajectory for MOE to track.
Figure 6: The MOE-Hair hardware system setup. MOE is attached to a 6 degrees of freedom (DoF) robot arm. A third-person RGBD camera tracks key points of the head and the participant’s hand. A calibration marker is used to transform between camera and robot coordinate frames.
@@ III-D MOE-Hair System Integration
Figure 7: The MOE-Hair system integration. We use a third-person RGBD camera to track the user’s visual face key points. The visual perception module maps the visual key points to the 3D point cloud and generates a task-specific trajectory. As MOE follows the trajectory, the system predicts the applied force on the head using the segmented depth images of the MOE fingers and actuator current load. The system adjusts the trajectories in real time to maintain comfort and contact based on the predicted contact forces.
@@ IV System Results
We present evaluation results for the components of the proposed MOE-Hair system. Specifically, we demonstrate that the proposed MOE soft robot end-effector can interact with a surface while applying lower contact forces compared to a rigid end-effector. Furthermore, we show that we achieve improved applied force estimation using both observed depth and actuator current loads, compared to using only depth or actuator load.
@@ IV-A MOE Interaction Force Evaluation
Figure 8: Visual comparison of the applied force estimation errors. We compare between the ground-truth forces experienced by the mannequin head with different wigs, to the predicted forces from the baselines and the proposed approach.
@@ IV-B Applied Force Estimation Evaluation
For three wigs with different lengths and color, we have MOE follow a finger combing trajectory and record the depth image and actuator current load observation along with the synchronized ground-truth force readings from the sensor attached to the mannequin head. We visualize the root mean squared error (RMSE) over the trajectory in Fig. 8 and report them in Table II. In all three cases, the proposed MOE-Hair applied force estimation module outperforms both baselines, improving the actuator load-only and depth image-only baselines by up to 60.1 % and 20.3 %, respectively.
@@ V User Study Evaluation
Figure 9: User study results. We evaluate on perceived task effectiveness, user comfort, and appropriate use of force. Top row: results for the head patting task where MOE pats the user’s head from the top and the side. Middle row: results for the finger combing task where MOE follows the user’s demonstrated trajectory to comb the hair. Bottom row: results for the hair grasping task where MOE approaches the user’s head and grasps with its dexterous soft fingers.
@@ V-B Participants
We recruited 12 participants (ages 20–27, hair lengths 25–508 mm) without visual or mobility limitations as this is the first work with user studies on robot haircare. The participants represented diverse hair types based on an accepted classification [55].
@@ V-D Responses
Fig. 9 shows the participant responses for task effectiveness, user comfort, and the system’s appropriate use of force in each of the tasks for both force feedback conditions. Additionally, we report participant responses for their preferred force feedback condition in Fig. 10, where higher Likert scale rating corresponds to a preference for MOE-Hair. We conducted non-parametric Wilcoxon signed-rank tests [57] for paired user responses based on the two systems performing the same tasks from Fig. 9. We also evaluated if the participants had a significant preference for a system from the neutral value of 4 with the responses in Fig. 10.
Task Effectiveness. Across all three tasks, participants rated the task effectiveness of MOE-Hair as higher than the vision-only baseline (all p<.01)p{<}.01). Additionally, we found a statistically significant preference for MOE-Hair for all three tasks for the comparison question of system effectiveness. The finding supports H1 that users will find MOE-Hair to be more effective for the tasks compared to the vision-only system.
User Comfort. The participants reported being comfortable with both systems in all of the tasks, generally agreeing that the system put them at ease and that they trusted the system to be safe. However, we only found significant differences in the hair-grasping task, where participants rated MOE-Hair as more comfortable than the vision-only baseline ($p=.02$). We found no statistically significant preference for either force feedback condition on user comfort, although the participants tended to be more comfortable with the MOE-Hair system. This partially supports H2 with the caveat that the biggest differences in comfort are for more involved tasks. The vision-only system sometimes failed to make consistent scalp contact, which some participants found similarly comfortable due to the lack of direct contact.
@@ V-E Qualitative Results
Effectiveness. All 12 of the participants responded positively to the system’s ability to perform the three skills outlined in this work. The participants indicated that MOE effectively performed the tasks and P6 noted, “It was working reliably.” P12 noted, “I didn’t think it was going to work on my braids but it actually did well.” Similarly, P2 remarked “I’ve been told my hair is difficult because it’s so thick” and commented that MOE-Hair performed effectively, noting “It did pretty well.” The responses underscored the system’s ability to perform well with varying hair characteristics and conditions.
Negative responses to the system’s effectiveness are related primarily to the participant’s uncertainty about the capabilities of the system. 3 of the participants responded that they wished the system could perform more tasks such as hair cutting and styling, where P11 commented “it would’ve been nice if it could cut my hair,” so that “I don’t have to talk to [human hairstylists].”
User tactile sensation. All 12 of the participants noted that they generally felt safe. For example, P7 noted that they “felt really comfortable.” 5 participants commented that MOE’s contact sensation was pleasant. P1 liked the sensation when MOE made contact, remarking “It was like a head massage.” P2 noted that MOE fingers felt “really similar to [human] fingers” and that “I kind of forgot it was a robot arm for a minute.” P11 stated that MOE “made me relax.”
Negative responses highlighted some participants’ hesitancy with an autonomous robot system touching their heads. P6 noted that MOE-Hair “felt like it’s kind of pushing my head.” P12 noted that they were at first uncertain if the system was “safe or not” with the first task but stated that they became “more comfortable” with the latter tasks, suggesting the process of MOE-Hair system building trust with its users.
Appearance of MOE. 8 participants made generally positive comments about MOE’s appearance. As P4 noted, “I felt safe because it looked rubbery and soft.” 3 participants noted that they liked that MOE did not look anthropomorphic and P5 stated MOE “just looks like it’s its own thing and it doesn’t give me an uncanny valley feeling.”
2 participants noted that the wires of the actuators contributed to their unease with the system. For example, P1 noted MOE “seemed like a prototype and doesn’t seem like a product” because of the visible wires. P7 noted MOE “doesn’t look terrible” but that “the cables could be cleaned up.” P12 made comments on MOE’s color and shape commenting “it doesn’t look friendly.” Diverging comments on the appearance and visual behavior of the robot could indicate a need to incorporate design insights from literature [58] to be more likely to be perceived as being friendly by the users.
@@ VI Discussion and Limitations
In this work, we introduce a soft robot manipulator for hair manipulation and care tasks that we call MOE. The results suggest that in comparison to its mechanically rigid counterparts, MOE is safer in close contact with a head and that observing deformations of MOE can provide us with useful information about the applied forces on human users. The experimental results with the mannequin testbed showed that soft robots such as MOE could be effectively exploited in hair care tasks. Human user study highlighted that the proposed MOE-Hair system with the contact force estimation module was perceived favorably in its effectiveness for the tasks, comfort, and ability to use appropriate forces. Qualitative analysis of the participants’ responses to open-ended questions revealed that some also perceived the tactile sensation of MOE to be pleasant and relaxing, which seems to offer promising extension in applying MOE to other contact-rich pHRI tasks.
@@
Another limitation is in the system’s behavior generation. The focus of the MOE-Hair system development was on effective and safe contact using the compliance of the MOE soft fingers and force feedback with the proposed applied force estimation module. However, the characteristics of movement and personalization of robot trajectories could be important factors for users viewing the system favorably [7]. To this end, we can improve MOE-Hair system to incorporate strategies for planning trajectories with human preferences. Future work will also explore customizing force thresholds and approach speeds for elderly and mobility-impaired users to maximize safety and comfort.
Acknowledgment
This work is supported by the NSF GRFP (Grant No. DGE2140739) and MOTIE, Korea (Grant No. 20018112).
@@ Appendix C Questionnaire
Appendix D Institutional Review
The Institutional Review Board approved the protocol for the trials and the study (STUDY2023_00000502).